OpenStudy (loser66):

What is wrong? From \(1/x -x) = (1/2x)+ (1/2)x\)

OpenStudy (loser66):

How to solve?

OpenStudy (wmj259):

\[\frac{ 1 }{ x }-x=\frac{ 1 }{ 2 }x+\frac{ 1 }{ 2 }x\] Is that the right equation?

OpenStudy (loser66):

yes

OpenStudy (johnweldon1993):

Oh it is? I thought that first fraction would be 1/(2x)

OpenStudy (loser66):

oh, @johnweldon1993 is correct. I am sorry @wmj259

OpenStudy (wmj259):

So the x is in the denominator.

OpenStudy (freckles):

so it is: \[\frac{1}{x}-x=\frac{1}{2x}+\frac{1}{2}x\] is this an equation you are suppose to be solving?

OpenStudy (wmj259):

@freckles, that looks correct.

OpenStudy (loser66):

yes, it is

OpenStudy (johnweldon1993):

So I get \(\large \pm \frac{1}{\sqrt{3}}\) solving for 'x'

OpenStudy (freckles):

did you try to get rid the fractions first multiply both sides by 2x

OpenStudy (johnweldon1993):

*I know no direct answers...but theres not much to do but simplify

OpenStudy (freckles):

and it is also @Loser66 I think she knows how to solve this...

OpenStudy (wmj259):

You cant multiply both sides by 2x since the fractions denominators aren't the same. I believe you first have to get a common denominator on the right side first.

OpenStudy (campbell_st):

1st statement should be x cannot equal zero

OpenStudy (loser66):

Feel free to give me the direct answer. hehehe...

OpenStudy (johnweldon1993):

Lol but it's easy to do \[\large \frac{1}{x} - x = \frac{1}{2x} + \frac{1}{2}x\] \[\large \frac{1}{x} - \frac{x^2}{x} = \frac{1}{2x} + \frac{x^2}{2x}\] \[\large \frac{1 - x^2}{x} = \frac{1 + x^2}{2x}\] \[\large 2x(1 - x^2) = x(1 + x^2)\] \[\large 2x - 2x^3 = x + x^3\] \[\large x = 3x^3\] \[\large 3x^3 - x = 0\] \[\large x(3x^2 - 1) = 0\] Then just \[\large 3x^2 - 1 = 0\] Unless I'm doing this wrong lol

OpenStudy (loser66):

Your solution is perfect. Thank you so much.

OpenStudy (wmj259):

I thought it was that the 2nd to last step, you can solve it so either x's = 0 so...... x=0 or x=+-sqrt(1/3)

OpenStudy (wmj259):

\[x=0 , x=\sqrt{1/3}\]

OpenStudy (johnweldon1993):

Well x cannot = 0 is the thing As you will not get an answer in the original problem *1/x, 1/(2x)

OpenStudy (wmj259):

Okay so plus or minus sqrt(1/3)